Arbitrage Rule of Thumb: If the difference in interest rates is greater than the forward premium/discount, or expected change in the spot rate for UIA, invest in the higher interest yielding currency. If the difference in interest rates is less than the forward premium (or expected change in the spot rate), invest in the lower yielding currency.

John Duell, a foreign exchange trader at JPMorgan Chase, can invest $8 million, or the foreign currency equivalent of the bank's short term funds, in a covered interest arbitrage with Denmark. Using the following quotes can John Duell make a covered interest arbitrage (CIA) profit? Use also the rule of thumb in answering the question.


Spot exchange rate Kr 6.172/$
3-month forward rate Kr 6.198/$
US dollar 3-month interest rate 3%
Danish kroner 3-month interest rate 4.5%
Answer #1
Premium/(discount) in the quotes = 6.198/6.172 -1 = 0.42%
Difference in interest rates = 4.5% - 3.0% = 1.50%
As the forward premium is not equal to the difference
interest rates, there is scope for CIA.
Since the difference in interest rates is more than the
forward premium, it would be profitable to borrow in the
currency having lower interest rate and investing the money
so borrowed in the currency having higher interest rate.
Steps to be taken:
1) Borrowing $8 million at 3%, the total amount payable
after 3 months being, 8000000*1.03 = 8240000 $
2) Convert the $8000000 to Kr at 6.198 to get 8240000*6.172 to get 50857280 Kr
3) Invest the Kr to get, after 1 year 50857280*1.045 = 53145858 Kr
4) Sell forward, Kr53145858 at 6.198 to get after 3 months
53145858/6.198 = 8574679 $
5) On the due date, realize the Kr deposit of 53145858, convert it to $8574679, repay the loan which has a maturity value of $8240000, the profit being 8574679-8240000 = 334679 $